# Neutral residues: Revisiting adapters for model extension

### v1

@@ 4.3 Main results
German
Transf-EN Backbone 43.0 60.3 40.5 56.5 34.5 33.8 30.2 24.0 44.3 27.8 47.0 32.0
finetuning 40.2 44.4 32.5 44.0 30.6 42.9 44.0 34.0 69.5 30.3 38.4 44.1
LoRA 43.4 52.0 36.5 50.6 32.6 40.8 40.9 30.4 64.0 29.9 43.0 41.2
Adapters 42.6 56.6 38.4 51.4 33.5 41.6 41.2 31.7 64.2 30.7 44.5 41.9
ours 42.9 59.5 40.5 56.3 34.3 39.7 41.5 31.2 62.3 30.5 46.7 41.0
Transf-ML Backbone 41.2 59.4 40.6 57.4 34.4 42.6 44.4 32.3 56.7 30.2 46.6 41.2
finetuning 42.1 50.7 37.8 51.5 32.6 42.3 47.5 35.6 70.2 31.8 42.9 45.5
LoRA 42.7 56.8 38.0 56.2 34.1 41.9 45.6 35.8 65.2 31.2 45.5 43.9
Adapters 41.2 58.0 39.0 55.1 34.6 39.3 43.3 32.3 56.0 29.7 45.6 40.1
ours 42.1 58.9 41.6 57.7 34.5 42.8 46.3 35.7 62.6 30.6 47.0 43.6
Gemma-2B Backbone 46.3 69.7 47.0 63.0 39.9 42.4 45.6 34.4 51.9 32.3 53.2 41.3
finetuning 45.1 55.1 40.5 54.4 34.3 44.9 50.1 37.4 68.8 33.1 45.9 46.8
LoRA 47.3 65.8 45.7 61.4 37.9 44.7 49.3 38.8 67.4 34.8 51.6 47.0
Adapters 45.9 65.9 47.4 60.9 39.5 44.7 46.9 36.5 63.6 34.5 51.9 45.2
ours 47.7 67.5 49.0 64.6 39.5 45.1 49.6 37.9 64.0 34.6 53.6 46.2
Table 6: Gating, activation and local loss. For the downstream tasks, we report the mean of all tasks.
Gate Gate losses Our Perplexity Tasks
$\ell_{1}$ CE init. En Fr En Fr
0.684 0.812 45.0 40.4
$\varnothing$ 0.681 0.813 45.5 40.6
0.697 0.812 45.6 41.0
0.668 0.810 45.7 41.1
Sigmoid 0.677 0.800 45.3 41.5
0.667 0.800 45.7 41.8
0.670 0.813 46.1 40.0
ELU 0.674 0.792 46.7 42.3
0.668 0.793 46.5 41.2
Backbone baseline 0.663 1.175 47.0 31.6
@@
Table 7: Ablation: $\bm{\alpha}$ governs the strength of the $\ell_{1}$ local loss. See Table 8 for more comparisons.
Perplexity Tasks avg.
$\alpha$ En Fr En Fr
$0$ 0.674 0.792 46.7 42.3
0001 0.674 0.792 46.4 42.0
0010 0.671 0.792 46.4 42.1
0100 0.668 0.793 46.5 41.2
1000 0.671 0.793 45.9 40.8
Table 8: Trade-offs with $\neq$ learning rates. This hyperparameter allows us to favor learning vs. forgetting or vice-versa. These results are the ones reported in Figure 1.
Perplexity Tasks avg.
Method LR En Fr En Fr
Backbone 0 0.663 1.175 47.0 31.6
$2\cdot 10^{-5}$ 0.678 0.963 45.3 36.3
Fine- $5\cdot 10^{-5}$ 0.705 0.840 35.9 42.1
tuning $2\cdot 10^{-4}$ 0.811 0.758 39.0 42.8
$5\cdot 10^{-4}$ 0.874 0.757 35.9 42.1
LoRA $2\cdot 10^{-5}$ 0.716 0.842 44.2 40.0
$5\cdot 10^{-5}$ 0.730 0.818 43.7 40.8
$2\cdot 10^{-4}$ 0.745 0.802 41.8 41.4
$5\cdot 10^{-4}$ 0.755 0.806 41.5 42.3
$2\cdot 10^{-5}$ 0.686 0.830 45.0 39.5
Vanilla $5\cdot 10^{-5}$ 0.687 0.812 45.0 40.4
adapters $2\cdot 10^{-4}$ 0.689 0.796 45.8 42.2
$5\cdot 10^{-4}$ 0.688 0.795 45.4 42.2
$2\cdot 10^{-5}$ 0.667 0.802 46.7 41.0
Neutral $5\cdot 10^{-5}$ 0.668 0.793 46.5 41.2
residues $2\cdot 10^{-4}$ 0.670 0.789 46.8 41.2
$5\cdot 10^{-4}$ 0.670 0.790 46.2 41.6
4.4 Ablations
Ablation of architectural components and losses. The low-variance initialization proposed in Section 3 improves the results in the context of gated adapters. Compared to He’s initialization, it favors training the gating faster that the other weights. This hypothesis is compatible with the observation that this initialization is detrimental without gating ($\varnothing$).
Ablation learning rate. Table 8 reports the trade-offs achievable by different techniques when varying the finetuning learning rates, see also Figure 1. The results of our method are remarkably stable across a large range of learning rate (one order of magnitude from $5\cdot 10^{-5}$ and $5\cdot 10^{-4}$), in sharp contrast with Fine-tuning and LoRA, which are highly sensitive to this parameter.
Ablation coefficient $\bm{\alpha}$. The perplexity results in Table 8 show that this weight gives a trade-off between learning and not forgetting, which is best optimized when selected $\alpha=100$. We retain this value in all our experiments. This choice does not translate to downstream tasks in this specific case, yet Table 8 shows that using the $\ell_{1}$ is preferable in other contexts (with no gating or sigmoid).
5 Conclusion
This paper has explored the feasibility and effectiveness of extending existing foundation models to incorporate new capabilities or knowledge without the need for resource-intensive retraining from scratch. By building upon the concept of adapters, which add new parameters to a pretrained model, we have demonstrated that it is possible to extend a model without compromising its original knowledge, thereby offering a more sustainable approach than retraining from scratch. Our study focused on the use-case of adding a new language to a pretrained model and evaluated the performance of the extended model on both training criteria and downstream tasks. The findings highlight several critical factors that contribute to the successful extension of a model while mitigating the issue of catastrophic forgetting. These factors include the strategic use of mixed training data, an adapter gating mechanism coupled with a local loss, and the importance of near-identical initialization.
Acknowledgments. This project is funded by Iliad Group, CMA CGM Group and Schmidt Sciences. We thank Xavier Niel, Rodolphe Saadé, Eric Schmidt, Aude Durand, Séverine Grégoire and Nicolas Granatino, for their support. We thank the Kyutai research team, with special appreciation to Patrick Pérez, Amélie Royer, Neil Zeghidour and Hippolyte Pilchen, for their valuable feedback, and Sarah Hôte and Guillaume Rouzaud for their help.
Appendix A architecture details and Hyper-parameter settings
Table 9: Pre-trained models: we consider two vanilla transformers that we have pretrained on two different corpus for which we know the data distribution: English only or MultiLingual with a majority of English. We also considered the Gemma 2B model (Gemma et al., 2024), for which we do not know the distribution. For this model, we refer to their paper for training hyper-parameters.
Transformer English Transformer Multilingual Gemma 2B
number of layers $L$ 16 16 18
working dimensionality 2048 2048 2048
number of heads 128 128 256
dimension FFN latent space 5632 5632 16384
activation type gated SILU gated SILU gated GELU
normalization RMS pre-normalization
group query
pretraining context length 4096 4096 8192
Hyper-parameters for training the extended model
batch size 64 64 8
context length 4096 4096 4096
#training steps 100000 100000 150000
AdamW: $\beta_{1}$ 0.9 0.9 0.9
AdamW: $\beta_{2}$ 0.95 0.95 0.95

### v3

@@ Appendix B Impact of the number of parameters
French
$p=0.1$ Backbone 43.0 60.3 40.5 56.5 34.5 34.6 32.8 26.7 42.3 28.0 47.0 32.9
Finetuning 37.9 45.5 34.8 45.5 30.9 42.2 48.6 35.3 60.3 31.5 38.9 43.6
LoRA 42.0 53.1 38.1 52.3 32.8 42.4 46.2 32.7 54.2 30.7 43.7 41.2
Adapters 43.1 57.7 40.3 54.1 34.5 44.3 48.8 34.8 52.7 31.2 46.0 42.4
ours 43.9 59.3 41.6 56.8 34.1 45.1 48.8 35.7 56.9 31.6 47.1 43.6
$p=0$ Backbone 43.0 60.3 40.5 56.5 34.5 34.6 32.8 26.7 42.3 28.0 47.0 32.9
Finetuning 36.4 42.1 32.5 43.0 29.4 42.4 49.2 36.4 60.3 32.0 36.7 44.1
LoRA 40.9 50.4 37.1 49.0 32.3 42.2 46.6 33.3 54.7 30.6 41.9 41.5
Adapters 42.2 55.1 38.4 52.8 32.9 42.3 48.3 33.6 53.5 31.5 44.3 41.9
ours 41.9 58.6 39.0 54.2 33.6 43.1 48.5 35.8 56.3 31.4 45.5 43.0
We also ran an experiment where EN-LM-1B was fine-tuned exclusively on Wikipedia data to assess the impact of this limited distribution on forgetting. Specifically, we fine-tuned EN-LM-1B on English Wikipedia for 50,000 steps using a learning rate of $2\cdot 10^{-4}$, as in Table 10, with the remaining hyperparameters listed in Table 8. This restricted fine-tuning dataset resulted in poor performance across all downstream tasks.
However, when combined with earlier results, this suggests that even when the fine-tuning distribution is only a rough approximation of the original pretraining distribution, strong performance on English downstream tasks can still be maintained.
Table 11: Impact of training on Wikipedia only.
Forgetting: English
Method BeleBele (0) HellaS (0) Arc C (5) CSQA (5) MMLU (5) Avg.
Backbone 43.0 60.3 40.5 56.5 34.5 47.0
Finetuning 39.0 49.3 36.7 47.6 31.0 40.7
Appendix D Math Dataset Experiments
We further fine-tuned the Gemma 2B model (Gemma et al., 2024) on two math-focused datasets—OpenMathInstruct-2 and MetaMathQA—for 50,000 steps using a batch size of 4096×644096$\times$ 644096 × 64 and a peak learning rate of $5\times 10^{-5}$ across all configurations. For the Ours variant, we set the L1 regularization coefficient $\alpha$ to 0.010.010.01. The results are presented in Table 12. We observe significant forgetting on English tasks, likely due to the overlap between the mathematical English in the fine-tuning data and general English. The Wikipedia-based English corpus is not sufficient to help the model disambiguate between the two.
Table 12: Maths results: We report the best method in bold, and the second best is underlined.
Forgetting: English Learning: Target Task avg.
Target Method BeleBele (0) HellaS (0) Arc C (5) CSQA (5) MMLU (5) GSM8K (5) Avg. score
Maths Backbone 47.1 71.3 47.8 64.5 39.0 20.0 53.9
Finetuning 43.6 64.1 35.1 56.4 31.8 58.4 46.2
LoRA 41.7 68.5 35.3 37.2 32.5 57.6 43.0
Adapters 42.8 68.6 38.1 58.7 32.4 47.6 48.1
Ours 42.9 68.2 39.2 60.8 32.0 54.4 48.6
Appendix E Japanese Dataset Experiments
To assess our approach beyond alphabetic scripts, we conducted additional experiments with Japanese, a language that uses a non-Latin script. Specifically, we augmented Gemma-2B with Japanese data extracted from CommonCrawl, processed using the same pipeline as the other multilingual finetuning datasets. The results are presented in Table 13.
Table 13: Results on Japanese: We report the best method in bold, and the second best is underlined.
Forgetting: English Learning: Target Task avg.
Target Method
BeleBele (0)
HellaS (0)
Arc C (5)
CSQA (5)
MMLU (5)
BeleBele (0)
CSQA (5)
MMLU (5)
Forgetting
Learning
Japanese Backbone 46.2 69.7 46.9 61.8 40.4 37.3 55.6 30.3 53.0 41.1
Finetuning 44.6 60.2 42.3 57.2 35.6 38.7 68.4 32.5 48.0 46.5
LoRA 45.7 63.3 45.5 58.7 37.4 37.3 63.0 32.4 50.1 44.2
Adapters 46.9 65.9 46.5 60.2 38.7 37.1 63.7 32.7 51.7 44.5
Ours 47.1 66.9 48.5 61.6 39.1 39.1 64.1 32.7 52.6 45.3
Appendix F Multilingual Dataset Experiments
We explore a multilingual setting where the goal is to learn new languages while preserving performance on previously known ones. We use the Gemma 2 2B model (Team et al., 2024), chosen for its strong multilingual capabilities. The learned languages are Danish, Hungarian, and Slovak; the retained ones are English, French, German, Portuguese, Spanish, and Italian. Training data consists of 90% balanced learned languages and 10% balanced retained languages ($p=0.1$), approximating the model’s original distribution. We compare all methods (with 20% additional parameters, except full fine-tuning) on their ability to learn new languages (Table 15) without degrading performance on retained ones (Table 14). We also evaluate them on FLORES (NLLB Team et al., 2024) (EN $\rightarrow$ X) in a 5-shot setting.
Table 14: Forgetting results for the multilingual setting: We report the best method in bold, and the second best is underlined.
Target Method BeleBele (0) HellaS (0) Arc C (5) CSQA (5) MMLU (5) FLORES (5) Avg.
English Backbone 56.3 73.9 65.6 69.1 52.3 x 63.4
Finetuning 46.1 59.2 40.9 53.1 33.7 x 46.6
LoRA 49.3 67.8 53.8 64.0 41.3 x 55.2
Adapters 52.2 72.8 64.8 68.1 50.4 x 61.7
Ours 52.8 73.3 65.4 66.6 51.6 x 61.9
French Backbone 48.8 60.1 56.2 58.5 44.9 43.9 52.1
Finetuning 45.4 50.5 37.6 42.7 30.4 34.7 40.2
LoRA 45.1 55.1 44.8 48.7 35.2 39.2 44.7
Adapters 47.6 58.7 55.4 59.3 42.6 43.6 51.2
Ours 48.8 59.7 57.1 59.9 44.0 44.1 52.3
German Backbone 47.6 54.0 52.5 65.0 43.7 30.8 48.9
Finetuning 42.4 47.5 35.3 50.8 30.8 25.2 38.7
LoRA 44.1 50.4 44.0 53.7 35.7 26.5 42.4
Adapters 47.2 52.7 53.5 65.8 42.7 30.6 48.8
Ours 47.2 53.1 54.4 68.4 43.9 30.6 49.6
Portuguese Backbone 46.3 58.7 56.5 68.3 45.1 44.3 53.2
Finetuning 40.9 50.8 37.1 51.6 30.7 37.2 41.4
LoRA 41.3 50.5 43.2 54.4 34.8 33.3 42.9
Adapters 43.9 57.4 55.3 69.5 43.1 44.5 52.3
Ours 44.4 58.1 56.7 68.6 44.4 44.0 52.7
Spanish Backbone 48.1 61.2 59.2 x 45.3 25.8 47.9
Finetuning 41.4 51.6 40.3 x 30.9 23.0 37.4
LoRA 43.8 56.1 45.2 x 36.6 24.7 41.3
Adapters 45.7 59.8 56.6 x 43.7 25.8 46.3
Ours 44.3 60.6 58.2 x 44.9 26.0 46.8
Italian Backbone 42.7 56.3 54.2 x 44.4 24.8 44.5
Finetuning 38.4 48.5 35.1 x 30.0 21.3 34.7
LoRA 39.8 48.2 41.1 x 34.8 19.4 36.7
Adapters 42.7 54.8 54.7 x 43.0 25.1 44.0
Ours 42.0 55.6 54.6 x 44.0 25.6 44.4
Table 15: Learning results for the multilingual setting: We report the best method in bold, and the second best is underlined.
Target Method BeleBele (0) HellaS (0) Arc C (5) MMLU (5) FLORES (5) Avg.
Danish Backbone 44.3 52.0 51.8 43.8 34.8 45.3
Finetuning 44.4 58.2 41.1 33.2 43.6 44.1
LoRA 44.6 60.4 46.9 37.1 44.2 46.6
Adapters 46.8 58.7 54.5 43.5 43.3 49.3
Ours 44.2 56.6 53.7 43.6 41.6 48.0
Hungarian Backbone 39.7 42.6 45.6 40.2 14.0 36.4
Finetuning 39.4 49.4 39.0 32.2 21.5 36.3
LoRA 33.4 35.0 33.2 29.3 20.8 30.3
Adapters 41.9 48.3 50.2 40.4 21.3 40.4
Ours 40.3 46.0 48.1 40.3 19.0 38.8
Slovak Backbone 43.3 46.4 49.7 41.6 19.2 40.1
Finetuning 43.9 50.7 38.8 33.1 28.6 39.0
LoRA 39.1 38.2 37.3 31.2 28.4 34.8
Adapters 44.9 50.5 50.7 41.6 28.1 43.1
Ours 43.7 48.6 49.5 41.6 25.8 41.8
Appendix G Detailed Ablation Results
This section provides the full detailed results corresponding to the ablation study discussed in section  4.4.
Table 16: Ablation: Gating and local loss. We report the best result in bold, and the second best is underlined.
Gate Local loss Forgetting: English Learning: French Task avg.
$\ell_{1}$ CE
BeleBele (0)
HellaS (0)
Arc C (5)
CSQA (5)
MMLU (5)
BeleBele (0)
HellaS (0)
Arc C (5)
CSQA (5)
MMLU (5)
Forget.
Learn.
(1) $\varnothing$ 43.1 57.4 37.5 54.5 33.7 43.2 47.7 34.6 54.9 31.7 45.3 42.4
(2) Sigmoid 41.0 57.9 40.2 53.1 34.0 43.2 47.4 36.2 54.5 31.4 45.2 42.6
(3) 42.3 57.7 39.1 53.7 33.4 44.0 47.6 34.5 52.3 31.1 45.3 41.9
(4) ReLU 43.4 59.1 39.3 55.8 34.4 44.6 48.7 34.8 54.3 31.8 46.4 42.8
(5) 43.9 59.3 41.6 56.8 34.1 45.1 48.8 35.7 56.9 31.6 47.1 43.6
(6) Adapter baseline 42.1 56.9 38.7 53.6 33.9 43.2 46.2 34.1 52.0 30.9 45.1 41.3
(7) Backbone baseline 43.0 60.3 40.5 56.5 34.5 34.6 32.8 26.7 42.3 28.0 47.0 32.9
Table 17: Ablation: Initialization. We report the best result in bold, and the second best is underlined.
Forgetting: English Learning: French Task avg.
Gate Init
BeleBele (0)
HellaS (0)
Arc C (5)
CSQA (5)
MMLU (5)
BeleBele (0)
HellaS (0)
Arc C (5)
CSQA (5)
MMLU (5)
Forget.
Learn.
Adapters w/ $\ell_{1}$-norm 43.9 57.2 39.2 53.6 33.5 42.4 46.6 34.0 53.6 31.3 45.5 41.6
43.1 57.4 37.5 54.5 33.7 43.2 47.7 34.6 54.9 31.7 45.3 42.4
Adapters w/ sigmoid 42.6 57.2 38.4 56.4 33.6 43.2 46.6 32.8 50.6 30.5 45.6 40.8
41.0 57.9 40.2 53.1 34.0 43.2 47.4 36.2 54.5 31.4 45.2 42.6
Neutral Residues 43.0 59.2 39.9 55.8 34.7 42.0 45.7 34.3 54.9 30.9 46.5 41.6
43.9 59.3 41.6 56.8 34.1 45.1 48.8 35.7 56.9 31.6 47.1 43.6
@@ Appendix G Detailed Ablation Results
Table 18: Ablation: $\boldsymbol{\alpha}$ controls the strength of the $\ell_{1}$ local loss. We report the best result in bold.
Forgetting: English Learning: French Task avg.
$\boldsymbol{\alpha}$
BeleBele (0)
HellaS (0)
Arc C (5)
CSQA (5)
MMLU (5)
BeleBele (0)
HellaS (0)
Arc C (5)
CSQA (5)
MMLU (5)
Forget.
Learn.
0 43.4 59.1 39.3 55.8 34.4 44.6 48.7 34.8 54.3 31.8 46.4 42.8
0.01 43.9 59.3 41.6 56.8 34.1 45.1 48.8 35.7 56.9 31.6 47.1 43.6
0.1 42.3 59.1 38.1 56.7 34.1 44.4 48.7 34.7 55.8 31.0 46.1 42.9
1 43.1 59.5 39.9 56.3 34.6 43.3 48.7 34.7 53.6 31.1 46.7 42.3
10 42.7 59.3 39.7 57.0 34.2 44.3 48.4 35.6 55.9 31.6 46.6 43.2
100 43.8 59.2 38.8 55.6 34.0 45.0 48.6 34.0 56.0 31.2 46.3 43.0
Table 19: Trade-offs with different learning rates. We report the best result in bold.
Forgetting: English Learning: French Task avg.
Method LR
BeleBele (0)
HellaS (0)
Arc C (5)
CSQA (5)
MMLU (5)
BeleBele (0)
HellaS (0)
Arc C (5)
CSQA (5)
MMLU (5)
Forget.
Learn.
Backbone 0 43.0 60.3 40.5 56.5 34.5 34.6 32.8 26.7 42.3 28.0 47.0 32.9
$2\cdot 10^{-5}$ 41.9 58.3 37.9 54.5 33.6 38.3 37.2 29.5 47.5 29.1 45.2 36.3
Fine-tuning $5\cdot 10^{-5}$ 41.8 55.8 37.5 53.7 33.1 41.7 44.1 34.0 54.8 30.8 44.4 41.1
$2\cdot 10^{-4}$ 37.9 45.5 34.8 45.5 30.9 42.2 48.6 35.3 60.3 31.5 38.9 43.6
$5\cdot 10^{-4}$ 36.9 40.4 32.1 41.1 29.5 42.8 47.6 34.0 57.3 30.9 36.0 42.5
LoRA $2\cdot 10^{-5}$ 42.3 54.7 38.9 51.5 32.9 40.9 44.4 32.4 54.7 30.1 44.1 40.5
$5\cdot 10^{-5}$ 42.0 53.1 38.1 52.3 32.8 42.4 46.2 32.7 54.2 30.7 43.7 41.2
$2\cdot 10^{-4}$ 41.3 51.1 36.1 48.4 32.4 42.1 47.3 33.8 53.9 31.2 41.9 41.7
$5\cdot 10^{-4}$ 41.1 50.8 36.6 47.0 32.3 44.7 48.0 34.4 54.8 31.0 41.6 42.6
Vanilla adapters $2\cdot 10^{-5}$ 43.0 56.9 38.7 52.8 34.0 41.7 44.6 33.3 54.3 30.3 45.1 40.8
$5\cdot 10^{-5}$ 42.1 56.9 38.7 53.6 33.9 43.2 46.2 34.1 52.0 30.9 45.1 41.3
$2\cdot 10^{-4}$ 43.1 57.7 40.3 54.1 34.5 44.3 48.8 34.8 52.7 31.2 46.0 42.4
$5\cdot 10^{-4}$ 42.3 57.9 39.2 53.8 33.6 45.2 48.6 36.2 56.2 31.5 45.4 43.5
Neutral residues $2\cdot 10^{-5}$ 42.6 59.1 40.8 55.5 33.9 44.9 47.4 35.6 54.6 31.9 46.4 42.9
$5\cdot 10^{-5}$ 43.9 59.3 41.6 56.8 34.1 45.1 48.8 35.7 56.9 31.6 47.1 43.6
$2\cdot 10^{-4}$ 42.7 59.1 40.7 55.0 34.5 43.6 49.2 35.2 53.1 31.6 46.4 42.5
$5\cdot 10^{-4}$ 42.8 59.3 39.4 57.2 34.2 44.0 49.1 35.2 55.2 31.2 46.6 43.0